OpenStudy (anonymous):

what are the vertex focus and directrix of the parabole with the giving equation y = 1/12 (x-4)^2-5

OpenStudy (campbell_st):

the standard vertex form is \[y = \frac{1}{4a}(x -h)^2 + k\] (h, k) is the vertex.... that's the 1st piece of information.... so match the equation to this to find h and k. the value a is called the focal length....(distance from the focus to the vertex) so compare \[\frac{1}{4a} = \frac{1}{12}\] and find the value of a. the focus will then be located at (h, k+a) the directrix equation is y = k - a hope it helps